OpenStudy (anonymous):

What's the correct simplification of the expression 4 over y to the power of negative 3.

OpenStudy (goformit100):

 is this the question ?

OpenStudy (anonymous):

 it looks like that

OpenStudy (goformit100):

Sorry, I don't know it :(

OpenStudy (anonymous):

ah its alright bud

OpenStudy (anonymous):

4 / y^-3 = 4y^3 Law of indices a / b^-n = ab^n

OpenStudy (anonymous):

Thank you